13) A robot stands on a large paved area. Beneath its feet is line that runs 39.96621 norh of east. As per its instructions it moves 11.29988° west of north for 10m. It then moved 4.61685° south of east until it reached the line. It then stopped. How far was the robot from its original position? (Hint: Tilt the coordinate system.)
Answer #1

The diagram for the given problem is shown below (making origin as initial position of the robot):

4.61685° 13.29988°10 m 39.96621 N :Northh S: South E East W: west

Now let's tilt the diagram by 11.299880 in clockwise direction. We will get this new diagram :

215.91673 10 m 28.666339 39.96621 - 11.29988 28.66633) (4.6168611.29988 15.91673)We have to calculate the value of OB . Here, OA = 10 m.

First we will derive equation of the line AB :

The coordinate of point A is (0 , 10).

The slope of the line AB can be calculated as

  mAB tan(360° 15.91673)tan(344.083270.2851732

So, the equation of line AB can be written as

y-10-一0.285 1732G-0)

or, y = 10-0.285 1732.r . ....................................(1)

Similarly, we can write equation of line OB.

The slope of OB will be

  m_{OB} = tan , (28.66633^{circ} ) = 0.54672046

As the line OB passes through the origin (0 , 0), we can write the equation of the line OB as

y-0 = 0.54672046(2-0

or, = 0.54672046 x ........................................(2)

Now, by solving equation (1) and (2), we can calculate the point of intersection of the line OB and AB or the coordinate of point B .

Putting the value of y from equation (2) in equation (1), wee get

0.54672046 r - 10 0.2851732 r

or,  0.83 1893 г. 10

or, x =12.020767 , , m

Putting the value of x in equation (2), we get

y =12.020767 imes 0.54672046 = , 6.571999 , , m

So, the coordinate of point B is (12.020767 , 6.571999) .

Now, we can calculate the length of the line OB as

OB V (12. 020767-0)2 + (6.571999-0)2 = V 144.49883927 + 43. 19117557

  = , sqrt{187.6900148} = , 13.700 , , m

Hence, the robot is 13.70 m far away from its original position.

For any doubt please comment and please give an up vote. Thank you .